OpenStudy (anonymous):

How do I determine the value of x that makes: x-4, 6, x, ... (arithmetic sequence)?

OpenStudy (asnaseer):

an arithmetic sequence is one where the difference between each term is the same. use that fact to come up with two equations and solve to find 'x'

OpenStudy (anonymous):

yup, I did but x-4?

OpenStudy (asnaseer):

steps are, difference between 1st and 2nd terms is:\[6-(x-4)=6-x+4=10-x\] then, difference between 2nd and 3rd terms is:\[x-6\] now, for an arithmetic sequence, these differences must be identical, therefore:\[10-x=x-6\] you should be able to solve this to find 'x'.